Find the measure of an angle where the compliment and three times the supplement add up to 242 degrees

Answer by ewatrrr(24785)

x the angle
(90-x) + 3(180-x) = 242
90-x + 540-3x = 242
630-242 = 4x
388 = 4x
97 = x check your numbers in question submitted